# SoK: Payment Channel Networks

### v2

@@ III Virtual channels
(b) Multi-hop virtual channel (the dots between Jill and Alice depict several intermediate nodes)
(c) Recursive virtual channel
@@
Protocol BC Validity Fee Privacy Off-chain dispute resolution Recursive Multihop
Perun [38] TC Limited
Jourenko et al. [40] UTXO Limited $\checkmark$
Aumayr et al. [41] UTXO Limited $\checkmark$ (fixed)
Elmo [42] UTXO Unlimited $\checkmark$
Donner [43] UTXO Unlimited $\checkmark$ (time based) $\checkmark$ $\checkmark$
Jia et al. [44] UTXO and TC Limited $\checkmark$ (fixed)
Eckey et al. [45] TC Limited $\checkmark$
Generalized state channels [46] TC Limited $\checkmark$
@@
Transactions in PCNs are routed from the sender to the receiver using a path of intermediate nodes. This pathfinding and routing mechanism requires the nodes along the payment path to be available for a transaction to be processed. Nodes, however, sometimes can choose to go offline or there can be network/service disconnections. This might cause the transaction to fail. Virtual channels, which are built on top of existing payment channels, provide a solution to this problem. Virtual channels facilitate payments between a pair of nodes using an intermediate node. The intermediate node has individual payment channels open with the other two nodes. The intermediary and the pair of nodes lock coins with each other in their respective payment channels to establish a virtual channel. Once the virtual channel has been established, the two nodes can be involved in a unlimited number of payments. These payments can be processed even if the intermediate node goes offline.
The first virtual channel construction, Perun, was proposed by Dziembowski et al. in [38]. Assume there are two users, Alice and Bob, along with an intermediary as depicted in Figure 3(a). Alice establishes a payment channel, $\alpha_{A}$, with the intermediary by locking $Y_{A}$ coins as collateral; the intermediary contributes $Y_{I}$ coins into this channel. Similarly, Bob locks $Z_{B}$ coins into his channel, $\alpha_{B}$, with the intermediary, which also deposits $Z_{I}$ coins from its possession. This marks the completion of the opening phase, effectively locking the coins belonging to Alice and Bob. Following this, Alice and Bob create a virtual channel to facilitate transactions between them. The virtual channel will include coins deducted from both Alice’s and Bob’s payment channels. Specifically, $X_{A}$ coins will be subtracted from $Y_{A}$ on Alice’s side and deposited into the virtual channel, while $X_{B}$ coins will be subtracted from $Z_{B}$ and will be deposited from Bob’s side. Now Alice and Bob can be involved in an unlimited number of payments without the intermediary being online.
@@
The idea of a virtual channel between a pair of nodes involving a single intermediary/intermediate node has been extended to establishing a virtual channel recursively over several hops involving several intermediaries, leading to the construction of a recursive virtual channel [40, 42, 43]. An example of a recursive channel is shown in the Figure 3(c). In the figure Trent and Alice are a pair of nodes who intend to establish a virtual channel between them. They do in three steps as follows; 1) Trent and Hu construct a virtual channel $V_{1}$ with Rajiv as an intermediary. 2) Hu and Alice construct a virtual channel $V_{2}$ with Jill as an intermediary. 3) Trent and Alice establish a virtual channel between them $V_{3}$.
@@
A multihop virtual channel is constructed by establishing a single virtual channel between a pair of users over a path comprising of several intermediate nodes. An example of multihop virtual channel is shown in the Figure 3(b). Rajiv and Alice establish a single virtual channel $V$ between them.
Motivation: We have already described how PCNs offer a solution to the blockchain scalability problem by facilitating instantaneous payments in an off-chain manner. PCNs however, have few limitations: 1) The path generated in a PCN to process a payment between a sender and a receiver usually contains one or several intermediate nodes. All the nodes along the path are required to be available to route the payment until the receiver. This may not be possible at all times since, nodes can be disconnected from the network due to various reasons such as power outages, DoS attacks, etc. In a virtual channel, once the coins of the intermediary(ies) are locked, it can choose to remain offline without affecting the transaction processing. 2) Each node along the path from the sender to the receiver charges a fees called as routing fees/forwarding fees for forwarding the payment to the next node along the path. The sender has to pay all the nodes their respective forwarding fee along the path. This increases the overall cost of routing the payment. Virtual channels provide a solution to this problem, since they do not charge fees on a per transaction basis. Instead, the fees is charged based on a time epoch basis (the time for which the coins of the intermediary/intermediaries are locked in the virtual channel) or a fixed fee can be agreed in advance by all the parties involved in the virtual channel construction. 3) The time taken to route a payment from a sender to a receiver is directly proportional to the path length. A longer path would result in an increase in the overall pathfinding and routing time. Virtual channels do not process payments through a path comprising of several intermediate nodes, hence the overall transactions’ pathfinding and routing time does not depend on the number of intermediate nodes between the sender and the receiver.
@@
Virtual channels should not add an unnecessary burden to users, and should mirror the operations of payment channels as closely as possible, with the added benefit of no on-chain transactions at all, while maintaining comparable security/privacy properties. We now give the properties desired from an ideal virtual channel, and compare the works in this area in Table II on the extent to which they achieve these properties.
Ideal properties:
1) BC: BC is the blockchain on which the corresponding virtual channel protocol can be deployed. TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO based scripting mechanism such as Bitcoin. Ideally, a virtual channel protocol should be deployable on any blockchain without the need for any additional programming/scripting requirements. The virtual channel protocol should not require any modifications to the existing consensus mechanisms of the underlying blockchains.
2) Validity: This determines the validity of the virutal channel. A limited validity means that the virtual channel is valid for a predetermined time period (which is decided by parties involved in the virtual channel). Unlimited validity means that the virtual channel can stay open until the parties involved decide to initiate closing. This provides better and efficient transaction processing between users who intend to have frequent payments made between each other.
3) Fee: This metric determines if the virtual channel takes into account the fees charged by the intermediate node(s) involved in the channel’s establishment. The fixed fee model implies that a predetermined, fixed fee is paid to the intermediate node(s) which is agreed upon by all the parties in the virtual channel and the time based fee model implies that a fixed fee is paid to the intermediary which depends on the time for which the virtual channel stays open. Having an well-defined fee structure will motivate the intermediary/intermediaries to participate in virtual channel creation.
4) Privacy: Similar to payment channels, we want privacy of nodes for virtual channels, which is said to be achieved when the identities of the nodes in the virtual channel protocols are not known to anyone except for their immediate neighbors.
5) Offchain dispute resolution: This metric determines if the transaction disputes in a virtual channel require a blockchain write.
6) Recursive: A virtual channel is said to be recursive or hierarchical if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing virtual channels on top of existing virtual channels from sender to receiver. We give an example of a recursive virtual channel in Figure 3(c). 7) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel between any pair of intermediate nodes along the path. We give an example of a multihop virtual channel in Figure 3(b).
@@
Protocol BC General/purpose state channels Security Graceful exit Off-chain dispute resolution
Pisa [47] TC $\checkmark$
Hydra [48] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Sprites [49] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Aumayr et al. [50] UTXO, TC $\checkmark$ $\checkmark$
Origami [51] TC $\checkmark$ $\checkmark$
Forcemove [52] TC $\checkmark$
@@ IV State channels
Motivation: State channels allow the already existing payment channels to facilitate execution of arbitrary applications such as games, e-commerce, etc. This leads to the PCN becoming more versatile. State channels should be deployable on any blockchain regardless of the underlying programming/scripting requirements and should be able to facilitate the execution of any decentralized application between the parties involved, while maintaining comparable security properties (in comparison to payment channels) and should ideally be able to resolve disputes without accessing the blockchain. We now give the properties desired from an ideal state channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of various state channel protocols in Table III. We describe the metrics used for comparison below.
Ideal properties:
1) Blockchain: This is the blockchain on which the corresponding state channel protocol can be deployed. TC is a blockchain that supports a Turing complete programming language and UTXO if a blockchain that supports a UTXO based scripting.
2) General purpose state channels: This is the ability of the state channel to facilitate the execution of any application in an off-chain manner by using a single application-agnostic instance of state channel on the blockchain.
3) Security: Security is achieved by a state channel protocol if no honest party loses funds because of malicious behavior by other parties in the protocol.
4) Graceful exit: A state channel protocol is said to employ graceful exit if it has clear and well-defined mechanisms for nodes joining or leaving the state channel.
5) Off-chain dispute resolution: This metric indicates if a dispute that occurs between the parties involved in the state channel can be resolved without accessing the blockchain.
@@ V Tumblers
TABLE IV: Comparison of Tumblers
@@
BOLT [53] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Nocust [54] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MIXCT [55] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
A2L [56]
UTXO and
TC
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Blindhub [57]
UTXO and
TC
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Teechain [58] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Accio [59] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Turbo [60] TC $\checkmark$ $\checkmark$
Boros [61] TC $\checkmark$ $\checkmark$ $\checkmark$
Tumblebit [62] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Motivation: A payment channel hub is a layer-2 mechanism that facilitates transaction processing between a sender/receiver pair that do not have a payment channel open between them. The intermediary which facilitates payments is called a tumbler. The tumbler has multiple payment channels open with several nodes in the network to facilitate payments. The hubs can potentially be useful in facilitating micropayments between a buyer and seller, without either of them needing to open a payment channel with each other.
Payment channel hubs can be classified into two types: on-chain and off-chain. Early hubs, also called hubs or mixers, were on-chain [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80], but they all suffered from scalability issues due to having to post each transaction on the blockchain. This led to the development of offchain payment channel hubs for specific blockchains, e.g., Bolt [53] for Zcash, Nocust [54] and MixCT [55] for Ethereum. The most general-purpose payment channel hubs are Tumblebit [62], $A^{2}L$ [56], and Blindhub [57].
@@
A payment channel hub should be able to facilitate payments between a sender and receiver, who do not have a payment channel open between them such that the hub cannot link a given transaction amount to a particular sender/receiver pair. Furthermore, the payment channel hub should also guarantee the fundamental property of atomicity (ensuring that the payment is sent to the receiver or it does not go through at all). We now give the properties desired from an ideal payment channel hub, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of off-chain payment channel hubs in the Table IV. We describe the comparison metrics below:
Ideal properties:
1) Blockchain: BC is the blockchain on which the corresponding Tumbler protocol can be deployed. TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO based scripting mechanism such as Bitcoin.
2) Relationship anonymity: It ensures that the relationship between a sender and a receiver for a given transaction should not be known to the tumbler.
3) Privacy against aborts: Tumbler should not be able to deduce the identities of a sender/receiver in case of a transaction abort, regardless of which party is responsible for the abort.
4) Time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch. A payment channel hub that supports transaction processing based on time epochs is not efficient because nodes cannot process payments between time epochs, which can be of arbitrary length. The duration between time epochs depends on several factors such as the unit of the time epoch (absolute time or relative time measured in terms of block height), number of users that will most likely involved in the next time epoch, etc. 5) Dynamic corruption: The payment channel hub should be able to handle the deviation of any party from the protocol at any point during its execution. i,e the set of adversaries should not be static during the protocol execution.
6) Atomicity: Atomicity ensured when a payment is routed all the way from the sender to the receiver or the payment is routed at all.
7) Value privacy: Value privacy is ensured by a payment channel hub, when, given a transaction amount, the tumbler cannot link it to a sender/receiver pair.
8) Variable amount: The tumbler should be able to process transactions of any value.
TABLE V: Comparison of Rebalancing Protocols
Protocol BC Trusted entity Graph compatability Privacy
Revive [81] TC $\checkmark$ (C) Cycles only
Hide & Seek [82] UTXO $\checkmark$ (D) Cycles only $\checkmark$
Subramanian et al. [83] TC and UTXO Agnostic $\checkmark$
Cycle [84] TC Cycles only $\checkmark$
Rebal [85] UTXO Cycles only $\checkmark$
Shaduf [86] TC Agnostic $\checkmark$
Chen [87] et al. UTXO Cycles
Musketeer [88] UTXO Cycles

### v3

@@ III Virtual channels
(b) Recursive virtual channel
(c) Multi-hop virtual channel (the dots between Jill and Alice depict several intermediate nodes)
@@
Protocol BC Validity Fee Privacy Off-chain dispute resolution Recursive Multihop Year
Generalized state channels [68] TC Limited $\checkmark$ 2018
Eckey et al. [69] TC Limited $\checkmark$ 2019
Perun [66] TC Limited $\checkmark$ 2019
Jourenko et al. [70] UTXO Limited 2020
Aumayr et al. [71] UTXO Limited $\checkmark$ (fixed) 2021
Elmo [72] UTXO Unlimited $\checkmark$ 2021
Donner [73] UTXO Unlimited $\checkmark$ (time based) $\checkmark$ $\checkmark$ 2023
Jia et al. [74] UTXO, TC Limited $\checkmark$ (fixed) 2023
@@
Motivation: Transactions in PCNs are routed from the sender to the receiver using a path of intermediate nodes. Current pathfinding and routing mechanisms require the nodes along the payment path to be available for a transaction to be processed. However, nodes can sometimes choose to go offline or there can be network/service disconnections causing transaction failures. Furthermore, each node along a payment path charges its own fees for forwarding the payment, which is paid by the sender and increases with the path length, hence the time taken to route a payment grows linearly in the path length. Virtual channels, which are built on top of existing payment channels, solve these problems. Initial constructions of virtual channels facilitated payments between a pair of nodes using a single intermediate node [66]. The intermediate node needs to have individual payment channels open with the other two nodes. The intermediary and the pair of nodes lock coins with each other in their respective payment channels and a virtual channel is established. Upon establishment of the virtual channel, the pair of nodes can be involved in a unlimited number of payments. These payments can be processed without the intermediate node being online. It might be better to use virtual channels from a routing fees perspective, since unlike routing protocols, nodes in virtual channels do not charge a routing fee for every transaction.
Alice, Bob, and an intermediary establish a virtual channel as shown in Figure 3(a). Alice locks $Y_{A}$ coins and the intermediary locks $Y_{I}$ coins in the payment channel $\alpha_{A}$. Similarly, Bob locks $Z_{B}$ coins and the intermediary locks $Z_{I}$ coins in their channel $\alpha_{B}$. The virtual channel $V$ is created once Alice locks $X_{A}$ coins from her balance in $\alpha_{A}$ and Bob locks $X_{B}$ coins from his balance in $\alpha_{B}$. Now Alice and Bob can process payments without the intermediary’s online presence.
@@
The idea of a virtual channel between a pair of nodes involving a single intermediary has been extended to establishing a virtual channel recursively over several hops involving several intermediaries, leading to the construction of a recursive virtual channel [70, 72, 73]. An example of a recursive channel is shown in Figure 3(b). A recursive virtual channel enables transaction processing between any pair of nodes, by recursively establishing virtual channels between several intermediaries, as opposed to [66], which facilitates virtual channels only between a pair of nodes connected directly to an intermediary.
@@
A multihop virtual channel is constructed by establishing a single virtual channel between a pair of users over a path comprising of several intermediate nodes. An example of multihop virtual channel is shown in Figure 3(c). A multihop virtual channel is an improvement over recursive virtual channels. In a multihop virtual channel, a single virtual channel can be established between any pair of nodes which are separated by several intermediaries.
@@
Virtual channels should not add an unnecessary burden to users, and should mirror the operations of payment channels as closely as possible, with the added benefit of no on-chain transactions at all, while maintaining comparable security/privacy properties. We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
Blockchain terminology: In the rest of the paper BC denotes a blockchain. TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO-based scripting mechanism such as Bitcoin. In Table II, we present a qualitative comparison of virtual channel protocols.
Ideal properties: 1) Validity: This determines the validity of the virtual channel. A limited validity means that the virtual channel is valid for a predetermined time period (which is decided by nodes involved in the virtual channel). Unlimited validity means that the virtual channel can stay open until the nodes involved decide to initiate closing [69]. Importance: Unlimited validity provides better and efficient transaction processing between users who intend to have frequent payments made between each other. 2) Fee: This metric determines if the virtual channel takes into account the fees charged by the intermediate node(s) involved in the channel’s establishment. The fixed fee model implies that a predetermined, fixed fee is paid to the intermediate node(s) which is agreed upon by all the nodes in the virtual channel before the channel establishment. The time-based fee model implies that the fee paid to the intermediary depends on the time for which the virtual channel stays open [69]. Importance: Having an well-defined fee structure will motivate the intermediary/intermediaries to participate in virtual channel creation. 3) Privacy: In any virtual channel construction, the real identity of a node should only be known to its immediate neighbor(s) [66]. Importance: Privacy is important since it helps in preserving topology privacy. 4) Offchain dispute resolution: This metric determines if the transaction disputes in a virtual channel require a blockchain write. Importance: Ideally, we would want a virtual channel construction to have off-chain dispute resolution since blockchain writes are expensive and time-consuming. 5) Support for multihop virtual channels with several intermediaries: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channels between any pair of intermediate nodes along the path. Importance: This property is ideal since it facilitates payments between any pair of nodes in the network by establishing only a single virtual channel, as opposed to a recursive virtual channel in which a sender/receiver, and the intermediate nodes lock coins in multiple virtual channels at the same time.
The most significant developments in virtual channels over the years are that newer protocols have incorporated a fees to be paid for the intermediary(ies) that lock coins in virtual channels and virtual channels now offer support for both TC based and UTXO based blockchains. Despite these developments, efficient virtual channel protocol design has three overarching research problems, which are discussed in Section VII.
@@
Protocol BC General/purpose state channels Security Graceful exit Off-chain dispute resolution Year
Forcemove [75] TC $\checkmark$ 2018
Pisa [76] TC $\checkmark$ 2019
Sprites [77] UTXO $\checkmark$ $\checkmark$ 2019
Hydra [78] UTXO $\checkmark$ $\checkmark$ 2021
Aumayr et al. [79] UTXO, TC $\checkmark$ 2021
Origami [80] TC $\checkmark$ $\checkmark$ 2023
@@ IV State channels
Motivation: State channels allow the already existing payment channels to facilitate execution of arbitrary applications such as games, e-commerce, etc. This leads to the PCN becoming more versatile. State channels should be deployable on any blockchain regardless of the underlying programming/scripting requirements and should be able to facilitate the execution of any decentralized application between the parties involved, while maintaining comparable security properties (to payment channels) and should ideally be able to resolve disputes without accessing the blockchain. We now give the properties desired from an ideal state channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of various state channel protocols in Table III. We describe the metrics used for comparison below.
Ideal properties: 1) General purpose state channels: This is the ability of the state channel to facilitate the execution of any application supported by the underlying blockchain in an off-chain manner Importance: This is an ideal feature since it facilitates the deployment of any application without accessing the blockchain, making state channels more versatile. 2) Graceful exit: A state channel protocol is said to employ a graceful exit if it has clear and well-defined mechanisms for nodes joining or leaving the state channel without accessing the blockchain. Importance: This is an important property since most of the state channel protocols require the user to interact with the blockchain when they join/leave the state channel(s) they are a part of. 3) Off-chain dispute resolution: This property indicates if a dispute that occurs between the parties involved in the state channel can be resolved without accessing the blockchain. Importance: Ideally, we would want a state channel construction to have off-chain dispute resolution since blockchain writes are expensive and time-consuming.
@@ V Tumblers
TABLE IV: Comparison of Payment Channel Hubs
@@
Year
BOLT [81] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2017
Tumblebit [82] UTXO $\checkmark$ $\checkmark$ 2017
Nocust [83] TC $\checkmark$ $\checkmark$ $\checkmark$ 2018
Teechain [84] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2019
$A^{2}L$ [85] UTXO, TC $\checkmark$ $\checkmark$ $\checkmark$ 2021
Accio [86] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2021
Boros [87] TC $\checkmark$ $\checkmark$ $\checkmark$ 2021
MIXCT [88] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2022
Turbo [89] TC $\checkmark$ 2022
Blindhub [90] UTXO, TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2023
Motivation: A payment channel hub (tumbler) is a multi-party off-chain system where users can establish payment channels with a central hub, which acts as an intermediary. It allows multiple users to send payments to each other without the need for direct payment channels between each user pair. The hub coordinates payments between different participants. The intermediary which facilitates payments is called a tumbler. Though a payment channel hub uses the same underlying infrastructure as that of PCNs and virtual channels, each of these constructions have their own use cases. PCNs are usually used when two nodes Alice and Bob transact on an infrequent basis. Virtual channels are used if Alice and Bob transact frequently, e.g., if Bob provides Alice with a service every month. Payment channel hubs are used when Alice needs to pay several receivers on a frequent basis and she does not want the tumbler to know the receivers.
@@
Payment channel hubs can be classified into two types: onchain and off-chain. Early hubs, also called hubs or mixers, were on-chain [91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102], but they all suffered from scalability issues due to having to post each transaction on the blockchain. The scalability issues of on-chain payment channel hubs have led to the development of offchain payment channel hubs for specific blockchains, e.g., Bolt [81] for Zcash, Nocust [83] and MixCT [88] for Ethereum. The most general-purpose payment channel hubs are Tumblebit [82], $A^{2}L$ [85], and Blindhub [90]. A payment channel hub should be able to facilitate payments between a sender and receiver, who do not have a payment channel open between them such that the hub cannot link a given transaction amount to a particular sender/receiver pair. Furthermore, the payment channel hub should also guarantee the fundamental property of atomicity (ensuring that the payment is sent to the receiver or it does not go through at all). We now give the properties desired from an ideal payment channel hub, and compare the works in this area on the extent to which they achieve these properties. In Table IV, we qualitatively compare several off-chain tumblers.
Ideal properties: 1) Relationship anonymity: It ensures that the relationship between a sender and a receiver for a given transaction should not be known to the tumbler [81]. Importance: This is an important property since one of the main uses and design goals of a tumbler is to facilitate anonymous transactions. Hence, any tumbler construction should satisfy this property. 2) Privacy against aborts: Tumbler should not be able to deduce the identities of a sender/receiver in case of a transaction abort, regardless of which party is responsible for the abort [90]. Importance: This is an important property to have since aborts can happen due to several reasons such as network disconnections, power outages, etc. Malicious nodes can also deliberately abort a protocol. In the case of tumbler protocols, the nodes whose payments did not succeed can be linked to each other once the protocol execution completes. 3) Independent of time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch [81]. This is not ideal. Importance: A payment channel hub that supports transaction processing based on time epochs is not efficient because nodes cannot process payments between time epochs, which can be of arbitrary length. The duration between time epochs depends on several factors such as the unit of the time epoch (absolute time or relative time measured in terms of block height), number of users that will most likely involved in the next time epoch, etc. 4) Protection against dynamic corruption of nodes: Current tumbler protocols corrupt nodes at the beginning of protocol execution and assume that the set of corrupted nodes remains constant until the protocol execution terminates. Ideally, a payment channel hub should be able to handle the deviation of any party from the protocol at any point during its execution. Importance: This is an important property since its infeasible from a practical standpoint to assume that the set of adversaries in the payment channel hub protocol remains constant during execution. 5) Atomicity: Atomicity is ensured if either the payment is routed all the way from the sender to the receiver or the payment is routed at all. Importance: This is important since it prevents honest parties from losing their funds because of malicious parties. 6) Value privacy: Value privacy is ensured by a payment channel hub, when, given a transaction amount, the tumbler cannot link it to a sender/receiver pair [81]. Importance: Value privacy is important since only the sender and receiver should know the amount being transacted. In the event of every sender/receiver pair transacting a unique amount, lack of value privacy will enable the tumbler to link a sender/receiver pair to a particular transaction, violating relationship anonymity. 7) Variable amount: The tumbler should be able to process transactions of any amount [81]. Importance: A tumbler having the ability to process amounts of variable value is important since its practically infeasible to assume that all sender/receiver pairs will always transact a fixed amount.
Over the years, various payment channel hub constructions have been developed to address specific challenges based on their unique design goals. However, we have identified three overarching research problems in payment channel hubs that require attention. We discuss them in Section VII.
TABLE V: Comparison of Rebalancing Protocols; C is centralized and D is distributed
Protocol BC Trusted entity Graph compatability Privacy Year
Revive [103] TC $\checkmark$ (C) Cycles only 2017
Subramanian et al. [104] TC and UTXO Agnostic $\checkmark$ 2019
Rebal [105] UTXO Cycles only $\checkmark$ 2021
Hide & Seek [106] UTXO $\checkmark$ (D) Cycles only $\checkmark$ 2022
Cycle [107] TC Cycles only $\checkmark$ 2022
Shaduf [108] TC Agnostic $\checkmark$ 2022
Musketeer [109] UTXO Cycles 2023
Chen [110] et al. UTXO Cycles 2024